OpenStudy (anonymous):

PLEASE HELP ME Write the equation in standard form of the line through J(4, 1)with slope –4.

OpenStudy (anonymous):

plug into the point slope formula y-y1=m(x-x1) :]

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

the standard form of a linear equation is y=mx+b

OpenStudy (anonymous):

YEAH

OpenStudy (anonymous):

so how can we get y-1=-4x+4 to look like that?

OpenStudy (anonymous):

wait i'm working it out now

OpenStudy (anonymous):

can you work it out and ill follow

OpenStudy (anonymous):

y=-4(4)+1= -15 Y=-4X-15?

OpenStudy (anonymous):

y-1=-4(x-4) = y-1=-4x+16 add 1 to both sides y=-4x+17

OpenStudy (anonymous):

and thats it in standard form ?

OpenStudy (anonymous):

yes, y=mx+b m is the slope of -4 and b is the y intercept of 17

OpenStudy (anonymous):

sorry i made a mistake on the math up there so i just deleted.

OpenStudy (anonymous):

thats what lilly21 said.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

wait so whats the right answer

OpenStudy (anonymous):

-17 MY BAD.

OpenStudy (anonymous):

y=-4x+17

OpenStudy (anonymous):

okay then i hope its right and thanks for the assistants

OpenStudy (anonymous):

good luck.

OpenStudy (anonymous):

im closing it